### v1

# Macroscopic suppression of supersonic quantum transport

Abstract
We consider a broad class of strongly interacting quantum lattice gases, including the Fermi-Hubbard and Bose-Hubbard models. We focus on macroscopic particle clusters of size $\theta N$, with $\theta\in(0,1)$ and $N$ the total particle number, and we study the quantum probability that such a cluster is transported across a distance $r$ within time $t$. Conventional effective light cone arguments yield a bound of the form $\exp(vt-r)$. We report a substantially stronger bound $\exp(\theta N(vt-r))$, which provides exponential suppression that scales with system size. Our result establishes a universal dynamical principle: macroscopic suppression of supersonic macroscopic transport (MASSMAT).
Lieb and Robinson [1] famously discovered that quantum lattice systems exhibit an “effective light cone” reminscient of relativistic systems. Their Lieb-Robinson bound (LRB) controls the probability that quantum information travels a distance $r>0$ in time $t>0$ by
$\exp(C(v_{\mathrm{LR}}t-r))$ (1)
for constants $C,v_{\mathrm{LR}}>0$. This establishes an effective light cone $v_{\mathrm{LR}}t=r$ beyond which information propagation is exponentially suppressed. The Lieb-Robinson velocity $v_{\mathrm{LR}}$ is an $\mathcal{O}(1)$ quantity proportional to the strength of local interactions.
As one of the few rigorous tools for analyzing strongly interacting quantum many-body systems, the LRB has proven remarkably powerful. Following breakthroughs of Hastings in the early 2000s [2, 3, 4], it was decisive in resolving fundamental problems across condensed matter physics, quantum information theory, and high-energy physics. Applications of LRBs include exponential clustering for gapped systems [2, 5], the definition and stability of topological quantum phases [2, 3, 6, 7, 8, 9, 10], the area law for the entanglement entropy [4], the control of dynamical entanglement generation [11, 12], the many-body adiabatic theorem [13, 14], quantum simulation algorithms [15, 16, 17, 18, 19], bounds on quantum messaging [20] and the fast scrambling conjecture [21, 22, 23, 24, 25]. Given the broad utility of LRBs, a large and continually growing body of research is concerned with extending therm and related propagation bounds to new settings, e.g., to long-range interactions [2, 5, 26, 27, 23, 12], open quantum systems [28, 29, 30], bosonic lattice gases [31, 32, 16, 33, 34, 35, 36, 37, 19, 38, 39], continuum systems [40, 41, 42]. Improved LRB establishing finer control (e.g., slow transport for disordered systems) have been proved in [43, 44, 45, 46, 47, 48, 9]. LRBs have also been observed experimentally [49, 50, 51, 52], e.g., with ultra-cold atoms in optical lattices. For a comprehensive review of progress up to 2023, see [53].
Ordinarily, it is considered a strength of the standard LRB (1) that it is independent of system size, making it well-suited for analyzing quantum dynamics on microscopic scales, where all relevant parameters ar $\mathcal{O}(1)$. However, many physically relevant problems concern the collective transport of macroscopic numbers of quantum particles, starting with Ohm’s law and ranging to the separation of timescales that is the basis of quantum hydrodynamics [54, 55, 56] and prethermalization phenomena [57, 58, 59]. Controlling macroscopic particle transport poses unique challenges — particularly in bosonic systems [60, 36, 61, 62, 63] which can exhibit large local particles numbers even within regions of $\mathcal{O}(1)$ size.
We establish a new type of dynamical bound on the transport of macroscopic particle clusters in strongly interacting quantum lattice systems. Specifically, we show that such transport is suppressed by an exceptionally rapid macroscopically large decay rate outside of the light cone: the bound takes the form
$\exp(CN(vt-r)),$ (2)
where $N$ is the total particle number. Figure 1 compares the standard light cone to this new MASSMAT light cone given by (2). What sets MASSMAT apart is the additional $N$-factor in the exponent in (2) compared to LRB-type bounds (1). Consequently, decay outside of the light cone $r>vt$ grows dramatically with increasing system size $N$. We refer to the bound (2) as a manifestation of a new quantum-dynamical principle: macroscopic suppression of supersonic macroscopic transport (MASSMAT). MASSMAT significantly strengthens prior bounds on the macroscopic particle transport problem [60, 36, 61, 62, 63].
A key conceptual consequence of MASSMAT is that the effective light cone $r=vt$ established by (2) becomes extremely sharp already for moderate $N$-values and mathematically exact (meaning free from errors) in the thermodynamic limit $N\to\infty$ — all while $r$ and $t$ are held fixed. This stands in contrast to the standard Lieb-Robinson light cone of the form (1), which is rougher because it allows for $\mathcal{O}(1)$ leakage. Hence, MASSMAT establishes that the transport of macroscopic particle cluster is universally governed by an unforeseen emergent strict causality. The well-known analogy connecting LRBs and special relativity through the shared concept of the light cone is thus shown to become an exact correspondence for thermodynamically large particle clusters thanks to MASSMAT. The precise, rigorous statement is given in Theorem 1 below.
Our proof of MASSMAT applies broadly to many strongly interacting quantum lattice gases, including the Fermi-Hubbard and Bose-Hubbard Hamiltonians with short-range hopping  [^1] In fact, the lattice structure is not essential either: what we use abstractly is the short range and bounded group velocity of the kinetic operator. These requirements also hold, e.g., for semi-relativistic electrons in the continuum which are described by the kinetic operator $\sqrt{-\Delta+m^{2}}$.. Therefore, MASSMAT is a universal dynamical principle that places unforeseen constraints on quantum many-body systems out of equilibrium.
$d_{XY}$time $t$$\mathcal{O}(1)$ decay rate(prior work)$\mathcal{O}(N)$ decay rate(MASSMAT)
Figure 1: Our main result establishes the green light cone $\sim vt$ with $v$ given by (6) outside of which the exponential decay rate becomes $\propto N$, i.e., macroscopically large. Since $v$ is larger than the quantity $\kappa$ in (9) that bounded the speed of macroscopic clusters in prior work [60, 36], there is a separation of the new macroscopic MASSMAT light cone and the usual $\mathcal{O}(1)$ light cone (yellow region).
Example: non-interacting chain. For illustration, we present a simple situation where MASSMAT obviously holds, while the decay provided by the LRB (1) is far too pessimistic. Consider the dynamics of a chain of free (i.e., non-interacting) bosons with only nearest-neighbor hopping, i.e., the Hamiltonian
$H_{\mathrm{free}}=\sum_{x=1}^{L-1}(a_{x}^{\dagger}a_{x+1}+a_{x+1}^{\dagger}a_{x}),$ (3)
For simplicity, consider the initial state where all particles are localized at site $1$, i.e., $\psi_{0}=(a^{\dagger}_{1})^{N}\Omega$ where $\Omega$ is the vacuum. We capture macroscopic transport via the projection $P_{N_{\{r,\ldots,L\}}\geq\theta N}$ onto the eigenspaces of $N_{\{r,\ldots,L\}}=\sum_{x=r}^{L}n_{x}$ (the number of particles sitting on sites $\{r,\ldots,L\}$) with eigenvalues at least $\theta N$. Using that the particles are non-interacting, one easily finds (see [65] for the details)
$\langle\psi_{t}|P_{N_{\{r,\ldots,L\}}\geq\theta N}|\psi_{t}\rangle\leq e^{NC(vt-r)}$ (4)
which proves that the MASSMAT principle holds for the non-interacting chain.
For non-interacting particles, the power $N$ in (4) arises because the particles are statistically independent. Of course, this argument breaks down completely for strongly interacting particles. Surprisingly, as we show, MASSMAT holds nonetheless. We are able to achieve this by devising a new way of deriving many-body propagation bounds that we coin geometric exponential tilting which departs from prior approaches to bounding macroscopic transport [60, 36, 61, 62, 63]. We explain the core idea of geometric exponential tilting after we presenting the main result.
Setup and main result. We consider a finite graph $(\Lambda,\mathcal{E}_{\Lambda})$ with vertex set $\Lambda\subset\mathbb{R}^{D},\,D\geq 1$, such that nearest-neighbors all have Euclidean distance $=1$.
We consider a system of indistinguishable quantum particles living on the graph $\Lambda$ — the result and its proof are identical for fermions and bosons and so we treat both cases in parallel. We take $\{a_{x}^{\dagger},a_{x}\}_{x\in\Lambda}$ to be a collection of fermionic/bosonic creation and annihilation operators satisfying the usual canonical anticommutation/commutation relations.
The model. On the fermionic/bosonic Fock space over the one-body Hilbert space $\ell^{2}(\Lambda)$, we consider Hamiltonians of the form
${H=H_{0}+V(\{n_{x}\}_{x\in\Lambda}),\quad H_{0}=\sum_{x,y\in\Lambda}J_{xy}a_{x}^{*}a_{y}}.$ (5)
Here, $J_{xy}$ represents short-ranged particle hopping and $V(\{n_{x}\}_{x\in\Lambda})$ represents a general density-density interaction.
Our assumptions on the Hamiltonian (5) are as follows:
  1. (i)
    The hopping matrix is Hermitian, i.e., $J_{xy}=\bar{J}_{yx}$ for $x,\,y\in\Lambda$, and satisfies, for a parameter $a>0$, the short-range condition
    $v=\max_{x\in\Lambda}\sum_{y\in\Lambda}|J_{xy}|\frac{\sinh(a|x-y|)}{a}<\infty.$ (6)
  2. (ii)
    $V:\{0,1,2,\ldots\}^{|\Lambda|}\to\mathbb{R}$ is a real-valued function of $|\Lambda|$ variables.
Under these assumptions, $H$ is a self-adjoint operator on the Fock space $\mathcal{F}(\ell^{2}(\Lambda))$; see, e.g., [35].
The class of Hamiltonians of the form (5) and satisfying these assumptions is very broad. In particular, it includes the paradigmatic Fermi-Hubbard and Bose-Hubbard Hamiltonians. We call Condition (i) the short-range condition because
$\sinh(a|x-y|)\sim\frac{1}{2}\exp(a|x-y|),\qquad|x-y|\gg 1$
and so (6) holds precisely when the hopping matrix $J_{xy}$ decays exponentially at large distances $|x-y|$. (We use the $\sinh$ in (6) instead of the exponential because it gives the asymptotically sharp value of $v$ in the limit $a\to 0$ as we explain after the main result.) In particular, Condition (i) holds for the physically most important case of nearest-neighbor hopping, i.e.,
$J_{xy}=J\delta_{|x-y|=1},$
in which case the the short-range condition (6) holds for any $a>0$ with $v=2dJ\tfrac{\sinh a}{a}$.
Finally, Assumption (ii) on $V$ is extremely weak. In particular, long-range and $k$-body interactions for any $k$ are allowed as long as they are of density-density type. Typical examples of $V(\{n_{x}\}_{x\in\Lambda})$ are polynomials. E.g., for the paradigmatic Bose-Hubbard Hamiltonian, one has
$V(\{n_{x}\}_{x\in\Lambda})=\sum_{x\in\Lambda}\left(n_{x}(n_{x}-1)-\mu n_{x}\right)$
We remark that it is easy to include local spin degrees of freedom in our setup and we only refrain from doing so to keep the notation simple. Spin degrees of freedom appear, e.g., in the standard Fermi-Hubbard Hamiltonian.
The main result. Since the Hamiltonian (5) preserves the total particle number $N_{\Lambda}=\sum_{x\in\Lambda}n_{x}$, we henceforth work on a fixed eigenspace of $N_{\Lambda}=N$ for a fixed $N\geq 1$. To state our main result, we introduce some notation. For any subset $S\subset\Lambda$, we define $S^{\mathrm{c}}=\Lambda\setminus S$ and
$N_{S}=\sum_{x\in S}n_{x},\qquad\bar{N}_{S}=\frac{N_{S}}{N}.$ (7)
For $0\leq c\leq 1$, we write $P_{\bar{N}_{S}\geq c},\,P_{\bar{N}_{S^{\mathrm{c}}}\leq c}$, etc. for the associated spectral projectors of $\bar{N}_{S}$. Finally, given two subsets of the lattice $X,Y\subset\Lambda$, we write $d_{XY}$ for their Euclidean distance.
Our main result is the following:
Theorem 1 (MASSMAT principle).
Consider a Hamiltonian $H$ of the form (5) satisfying Assumptions (i)(ii) with $v,\,a>0$.
Then, for any $0\leq\alpha<\beta\leq 1$, and any disjoint subsets $X,\,Y\subset\Lambda$, the following estimate holds on each $N$-particle sector:
$\|P_{\bar{N}_{X}\geq\beta}\,e^{-\mathrm{i}tH}P_{\bar{N}_{Y}\geq 1-\alpha}\|\leq\,e^{-aN((\beta-\alpha)d_{XY}-v|t|)}.$ (8)
This theorem is proved in [65]. The bound (8) implies the following strong light cone estimate on the quantum probability that a macroscopic cluster comprised of $(\beta-\alpha)N$ particles traverses the distance $d_{XY}$ in time $t$. Indeed, consider two disjoint regions $X,Y\subset\Lambda$ and an initial $N$-particle quantum state $\psi_{0}$ that has at least $(1-\alpha)N$ particles in $Y$ and thus satisfies $\psi_{0}=P_{\bar{N}_{X}\geq 1-\alpha}\psi_{0}$. Consequently, there are at most $\alpha N\ll\beta N$ particles in $X$ initially and so $P_{\bar{N}_{Y}\geq\beta}\psi_{0}=0$. We denote the time-evolved state by $\psi_{t}=e^{-\mathrm{i}tH}\psi_{0}$. Then $\langle\psi_{t},P_{\bar{N}_{Y}\geq\beta}\psi_{t}\rangle$ is the probability that after time $t$, at least $(\beta-\alpha)N$ particles were transported from region $Y$ to the region of interest $X$. Thanks to (8), this probability is bounded by
$\langle\psi_{t},P_{\bar{N}_{X}\geq\beta}\psi_{t}\rangle=$ $\|P_{\bar{N}_{X}\geq\beta}\psi_{t}\|^{2}$
$=$ $\|P_{\bar{N}_{X}\geq\beta}e^{-\mathrm{i}tH}P_{\bar{N}_{Y^{\mathrm{c}}}}\psi_{0}\|^{2}$
$\leq$ $e^{-2aN((\beta-\alpha)d_{XY}-v|t|)}.$
In words, (8) implies that a macroscopic cluster of $(\alpha-\beta)N$ particles move at most at speed $\frac{v}{\alpha-\beta}$ up to errors that are exponentially small in $N$ and thus effectively completely negligible.
A few remarks about the bound (8) are in order. First, since it is an operator norm bound, it provides state-independent constants. Second, in applications, we do not need to require that $\psi_{0}=P_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha}\psi_{0}$ holds exactly and we can also consider mixed initial states. Third, as mentioned above, the propagation speed (i.e., the slope of the MASSMAT light cone) is given by $\frac{v}{\beta-\alpha}$. The constant $v$ is related to previous velocity bounds on particle transport [60, 36] as follows. Since $\frac{\sinh z}{z}\geq 1$ for $z\geq 0$, we have that Assumption (6) implies
$\kappa=\max_{x\in\Lambda}\sum_{y\in\Lambda}|J_{xy}||x-y|<v.$ (9)
This $\kappa$ is exactly the first moment of the hopping matrix which was used to bound the propagation speed in our prior works [60, 36]. Thus (9) shows that the maximal velocity / light cone slope $v$ for our light cone here is slightly larger than the slope $\kappa$ obtained in [60, 36]. This shows that the macroscopic decay rate outside of the MASSMAT light cone (which has slope $v$) sets on slightly later than the standard $\mathcal{O}(1)$ decay (compare (1)) that was proved in [60, 36]. This is shown in Figure 1. In fact, our choice of $v$ in Condition (i) is sharp in the long-range limit $a\to 0$: Using $\sinh(a|x-y|)\sim a|x-y|$ we see that $v$ in (6) converges to $\kappa$ as $a\to 0$. Compared to [60, 36], we see that by slightly increasing the light cone slope, we are able to boost the microscopic error estimate outside of the light cone to an unprecedented, macroscopic one.
Description of the proof method. Our proof of Theorem 1 rests on a approach which we call geometric exponential tilting. Here, we give a high-level overview and compare the approach to other ones in the literature. The full proof is deferred to [65]. Geometric exponential tilting is completely different to the approaches used in prior works to bound transport of macroscopic boson clusters, namely the second-order adiabatic spacetime localization observables (ASTLO) method [60, 36] and the optimal transport method [62]. The overarching idea of geometric exponential tilting is simple: We introduce a suitably chosen, invertible (but not unitary) many-body similarity transformation $T$ and then we bound the left-hand side of (8) by
$\|P_{\bar{N}_{Y}\geq\beta}\,e^{-\mathrm{i}tH}P_{\bar{N}_{X}\geq 1-\alpha}\|$ (10)
$\leq$ $\|P_{\bar{N}_{Y}\geq\beta}T^{-1}\|\|Te^{-\mathrm{i}tH}T^{-1}\|TP_{\bar{N}_{X}\geq 1-\alpha}\|.$
The first and third norm will produce the spatial decay $e^{-aN(\beta-\alpha)d_{XY}}$. The middle norm of $Te^{-\mathrm{i}tH}T^{-1}$ (which we call the deformed propagator) will produce the growth in time $e^{aNv|t|}$ for $t\in\mathbb{R}$, and so (8) follows.
The crux, of course, lies in choosing the right similarity transformation $T$. We construct a $T$ that exponentially weights the local particle numbers in a site-dependent way, i.e., $T=\exp(\sum_{x}F(x)n_{x})$ for a suitable real-valued function $F(x)$. The function $F(x)$ interpolates space-adiabatically between being $1$ on the region $X$ and $=-1$ on the region $Y$. Thus, $T$ gives large weight to configurations with many particles in $X$ and small weight to configurations with many particles in $Y$. The bound $\|Te^{-\mathrm{i}tH}T^{-1}\|\leq e^{aNv|t|}$ shows that the exponential weights in $T$ grow at most linearly in time under the dynamics.
The geometric exponential tilting method is simultaneously conceptually simple (recall (10)), flexible (one can adapt the similarity transform $T$ to the problem at hand) and powerful (it is so far the only method that yields MASSMAT).
The method has various links to prior works. First, it is inspired by a recent complex analysis argument for deriving transport bounds on non-interacting particles in [66]; see also [67]. The connection to complex analysis arises through Paley-Wiener theory [68], which in particular says that it is equivalent to have exponential decay in position space and to have an analytic extension of the Fourier transform to a complex strip  [^2] Indeed, Condition (i) is equivalent to saying that the family of transformed Hamiltonians $H_{\xi}=T_{\xi}HT_{\xi}^{-1}$, where $T_{\xi}=e^{\operatorname{\mathrm{d}\Gamma}(\xi\cdot x)}$, has an analytic continuation, $H_{\zeta}$, into the strip $\mathcal{S}_{a}^{n}=\left\{\zeta\in\mathbb{C}^{n}:\left\lvert{\rm Im}\zeta_{j}\right\rvert<a\;\forall j\right\}$. This analytic continuation breaks the time-reversal symmetry and one takes ${\rm Im}\zeta_{j}>0$ or ${\rm Im}\zeta_{j}<0$ $\forall j$, depending on whether one considers $t>0$ or $t<0$.. From a broader perspective, using suitable similarity transforms with locally varying exponential weights to adapt the geometry to the question at hand has a long history in mathematical physics, perhaps most famously in Witten’s proof of the Morse inequalities [70]. In the context of propagation bounds on quantum many-body systems, related uses of spatially varying weights have recently appeared in [34], Osborne-Yin-Lucas [71], and Fresta-Porta-Schlein [72] for different quantum-dynamical problems.
Conclusions. In this work, we have identified and rigorously proven a conceptually novel, universal bound on the nonequilibrium dynamics of strongly interacting quantum lattice models: the macroscopic suppression of supersonic macroscopic transport (MASSMAT). This principle establishes that the quantum probability of supersonic propagation of macroscopic particle numbers actually decays exponentially at a macroscopic rate proportional to the total particle number $N$. This is in stark contrast to what one obtains from Lieb-Robinson bounds which give an $\mathcal{O}(1)$ decay rate that does not grow with $N$. MASSMAT substantially strengthens the decay rate achieved on macroscopic boson transport in prior works [60, 36, 62].
The MASSMAT principle is universal in scope: It applies to both bosons and fermions (as well as mixtures) and holds across general geometries. Our proof is based on a new analytical technique — geometric exponential tilting — that is inspired by complex analysis methods from one-body quantum mechanics and developed here for the first time in a many-body context. We anticipate that this method will find broader applications in macroscopic transport problems, especially in regimes characterized by slow transport of large clusters, such as hydrodynamic limits [54, 55, 56] or prethermalization phenomena [57, 58, 59].
Our work opens several avenues for future exploration. One key question is the experimental observation of MASSMAT, e.g., in ultracold quantum gases on optical lattices. This requires observation of particle numbers that are large enough so that the improved decay outside of the MASSMAT light cone becomes observable. Another important avenue is to investigate if the MASSMAT principle extends to systems with long-range hopping, as studied in [60, 36, 61, 62, 63].
Data availability
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Acknowledgments
The research of J.F. is supported by the ANR, project ANR-22-CE92-0013. The research of M.L. is supported by the DFG through the grant TRR 352 – Project-ID 470903074 and by the European Union (ERC Starting Grant MathQuantProp, Grant Agreement 101163620). I.M.S. is supported by NSERC Grant NA7901. J.Z. is supported by National Key R & D Program of China Grant 2022YFA100740, China Postdoctoral Science Foundation Grant 2024T170453, National Natural Science Foundation of China Grant 12401602, and the Shuimu Scholar program of Tsinghua University.
Supplemental Material:
Macroscopic suppression of supersonic quantum transport
Jérémy Faupin, Marius Lemm, Israel Michael Sigal, and Jingxuan Zhang
This appendix has two parts. In Part I, we give the short proof of MASSMAT for the special case of a chain of non-interacting bosons (3), which was displayed in the main text as (4). In Part II, we introduce the geometric exponential tilting method and give the full proof of our main result, Theorem 1.
I Direct proof of MASSMAT for non-interacting bosons
In this appendix, we prove that (4) holds for the non-interacting Hamiltonian (3) by a short calculation. We consider the more general initial state $\psi_{0}=(a^{\dagger}(f))^{N}\Omega$ where $\Omega$ is the vacuum and $f:\{1,\ldots,L\}\to\mathbb{C}$ is a one-body wave function which is localized around the origin. Since the particles are non-interacting,
$\psi_{t}=e^{-\mathrm{i}tH}\psi_{0}=(a^{\dagger}(e^{-\mathrm{i}t\Delta_{L}}f))^{N}\Omega$
and so
$\langle\psi_{t}|P_{N_{\{r,\ldots,L\}}\geq\theta N}|\psi_{t}\rangle$ (S1)
$=$ $\sum_{N^{\prime}=\lceil\theta N\rceil}\left(\sum_{x=r}^{L}|\langle e^{-\mathrm{i}t\Delta_{L}}f,\delta_{x}\rangle|^{2}\right)^{N^{\prime}}.$
For the one-body Laplacian, it is easy to check from Fourier theory that $|\langle e^{-\mathrm{i}t\Delta_{L}}f,\delta_{x}\rangle|^{2}\leq e^{C(vt-r)}$ for suitable constants $C,v^{\prime}>0$. Therefore,
$\langle\psi_{t}|P_{N_{\{r,\ldots,L\}}\geq\theta N}|\psi_{t}\rangle\leq e^{NC(v^{\prime}t-r)}$ (S2)
The argument can be easily adapted to include on-site external fields, i.e., to treat Hamiltonians of the form
$H_{\mathrm{free}}=\sum_{x=1}^{L-1}(a_{x}^{\dagger}a_{x+1}+a_{x+1}^{\dagger}a_{x}+v_{x}n_{x}).$
with $v_{x}$ a bounded sequence. Indeed, one has
$\langle\psi_{t}|P_{N_{\{r,\ldots,L\}}\geq\theta N}|\psi_{t}\rangle$ (S3)
$=$ $\sum_{N^{\prime}=\lceil\theta N\rceil}\left(\sum_{x=r}^{L}|\langle e^{-\mathrm{i}t(\Delta_{L}+V_{L})}f,\delta_{x}\rangle|^{2}\right)^{N^{\prime}},$
with $V_{L}(x)=v_{x}$ for $1\leq x\leq L$. The one-body propagation bound of the form $|\langle e^{-\mathrm{i}t(\Delta_{L}+V)}f,\delta_{x}\rangle|^{2}\leq e^{C(vt-r)}$ follows, e.g., from [67, 66].
II Proof of Theorem 1
The proof of Theorem 1 is organized as follows.
  • In Step 1, we render the relative geometry of two disjoint subsets $X$ and $Y$ effectively one-dimensional by constructing a “separation function” $s(x)$ that incorporates the relevant geometry.
  • In Step 2, we introduce the exponential tilting operator $T$, which involves a similarity transformation that exponentially weighs the local particle numbers in a site-dependent way. The relative geometry between $X$ and $Y$ is fully taken into account through the function $s(x)$ defined in (S4), resp. (S7).
  • In Step 3, we derive the spatial decay from the first and third terms in (10).
  • In Step 4, we bound the tilted deformed propagator, i.e., the middle term in (10) and conclude the proof.
Step 1: Separating functions
Simple geometry: disjoint convex hulls. To fix ideas, we first consider the simplified scenario of two subsets $X,\,Y\subset\Lambda$ whose convex hulls, $\tilde{X}=\operatorname{conv}(X)$ and $\tilde{Y}=\operatorname{conv}(Y)$, are disjoint. With a slight modification of the argument, the results extend to complete general disjoint subsets; see the construction (S6) and (S7).
Let $x_{0}\in{\tilde{X}}$, $y_{0}\in{\tilde{Y}}$ be such that $d_{{\tilde{X}}{\tilde{Y}}}=|x_{0}-y_{0}|$ and introduce the “center of mass” coordinates
$w_{0}=\frac{1}{2}(x_{0}+y_{0}),\qquad b=\frac{x_{0}-y_{0}}{|x_{0}-y_{0}|}.$
By construction, the hyperplane $\left\{z\in\mathbb{R}^{D}:(z-w_{0})\cdot b=0\right\}$ separates ${\tilde{X}}$ and ${\tilde{Y}}$ to two different sides. Below we project the relative geometry of ${\tilde{X}}$ and ${\tilde{Y}}$ onto the line joining the points $x_{0}$ and $y_{0}$.
We introduce the separating function $s:\mathbb{R}^{D}\to\mathbb{R}$,
$s(x)=b\cdot(x-w_{0})$ (S4)
and define the following subsets of $\Lambda$ (see Figure S1)
$Y^{\infty}=\big{\{}x\in\Lambda\,|\,s(x)\leq-\frac{1}{2}d_{{\tilde{X}}{\tilde{Y}}}\big{\}},$ (S5)
$W^{0}=\big{\{}x\in\Lambda\,|\,-\frac{1}{2}d_{{\tilde{X}}{\tilde{Y}}}<s(x)<\frac{1}{2}d_{{\tilde{X}}{\tilde{Y}}}\big{\}},$
$X^{\infty}=\big{\{}x\in\Lambda\,|\,\frac{1}{2}d_{{\tilde{X}}{\tilde{Y}}}\leq s(x)\big{\}}.$
[…]
Figure S1: Schematic diagram for the decomposition (S5).
For the simple geometry, we have the following easy lemma.
Lemma 2.
Assuming $X$, $Y$ have disjoint convex hulls, we have $X\subset\tilde{X}^{\infty}$ and $\tilde{Y}\subset Y^{\infty}$.
Proof.
Let $x\in\tilde{X}$. We have
$s(x)=$ $b\cdot(x-x_{0})+b\cdot(x_{0}-w_{0})$
$=$ $\frac{x_{0}-y_{0}}{|x_{0}-y_{0}|}\cdot(x-x_{0})+\frac{1}{2}|x_{0}-y_{0}|.$
The second term equals $\frac{1}{2}d_{{\tilde{X}}{\tilde{Y}}}$ by the choice of $x_{0},\,y_{0}$, while the first term is non-negative by the separating plane theorem for disjoint convex sets [73]. This shows that $s(x)\geq\frac{1}{2}d_{\tilde{X}\tilde{Y}}$ and hence that $\tilde{X}\subset X^{\infty}$. The proof that $\tilde{Y}\subset Y^{\infty}$ is analogous. ∎
Extension to general geometry. Consider now arbitrary disjoint subsets $X,Y\subset\Lambda$. Let $g(x)=\operatorname{dist}_{Y}(x)-\operatorname{dist}_{X}(x)$. The separating hyperplane is now replaced by the separating hypersurface
$S=\left\{g(x)=0\right\},\quad\Omega_{\pm}=\left\{\pm g(x)>0\right\}.$ (S6)
Indeed, the hypersurface $S$ is equidistant to $X$ and $Y$, with $X\subset\Omega_{+}$, $Y\subset\Omega_{-}$; see Figure S2.
[…]
Figure S2: Schematic diagram for $S$ and $\Omega_{\pm}$.
In this more general case, we take the separating function $s(x)$ to be the signed distance function to $S$ with the sign chosen such that ±$\pm s(x)>0$ on $\Omega_{\pm}$. Explicitly,
$s(x)=\operatorname{sgn}(g(x))\operatorname{dist}_{S}(x).$ (S7)
(This reduces to (S4) in the simplified scenario considered before.) We note for later reference that this function is Lipschitz continuous with constant $1$, i.e.,
$\left\lvert s(x)-s(y)\right\rvert\leq\left\lvert x-y\right\rvert,\quad x,\,y\in\Lambda.$ (S8)
Indeed, on the same side of $S$, the function $s(x)$ coincides with the distance function up to a sign, which is 1-Lipschitz by the following standard argument. For any $z\in S$ and $x,\,y\in\Lambda$, we have $\operatorname{dist}_{S}(x)\leq|x-z|\leq|x-y|+|y-z|$ for any $S$. By taking $\inf_{z\in S}$, we obtain $\operatorname{dist}_{S}(x)\leq|x-y|+\operatorname{dist}_{S}(y)$. If $x,y$ fall on different sides of $S$, then we join them with a line segment passing $S$ at, say, $z$, and then apply the triangle inequality to $\operatorname{dist}_{S}(x)\leq|x-z|$ to conclude.
Similarly to (S5), we decompose $\Lambda$ with $s(x)$ from (S7) as follows:
$Y^{\infty}=\big{\{}x\in\Lambda\,|\,s(x)\leq-\frac{1}{2}d_{{X}{Y}}\big{\}},$ (S9)
$W^{0}=\big{\{}x\in\Lambda\,|\,-\frac{1}{2}d_{{X}{Y}}<s(x)<\frac{1}{2}d_{{X}{Y}}\big{\}},$
$X^{\infty}=\big{\{}x\in\Lambda\,|\,\frac{1}{2}d_{{X}{Y}}\leq s(x)\big{\}},$
As in Lemma 2, we have
Lemma 3.
We have $X\subset X^{\infty}$ and $Y\subset Y^{\infty}$.
Proof.
Consider any $x\in X$. On the one hand, we have $s(x)=\operatorname{dist}_{S}(x)$ by Definition (S7) and the fact that $X\subset\Omega_{+}$. On the other hand, since $S$ is equidistant to $X$ and $Y$, we have $d_{SX}=\frac{1}{2}d_{XY}$. This shows that $s(x)\geq d_{SX}=\frac{1}{2}d_{XY}$ and hence that $X\subset X^{\infty}$. The proof of $Y\subset Y^{\infty}$ is analogous. ∎
Step 2: Exponential tilting operator
In this section, we introduce the exponential tilting operator $T$; see (S16) below.
For brevity, we fix disjoint $X,\,Y\subset\Lambda$ throughout and denote
$d=d_{XY}.$ (S10)
We define the function $f:\mathbb{R}\to\mathbb{R}$ by
$f(s)={\bf 1}_{s\geq 1/2}-{\bf 1}_{s\leq-1/2}+2s{\bf 1}_{\left\lvert s\right\rvert<1/2}$ (S11)
as shown in Figure S3. (Here, ${\bf 1}_{\ldots}$ is the indicator function which equals $1$ if condition $\ldots$ is satisfied and which equals zero otherwise.)
$s$$f(s)$$\left(-\frac{1}{2},-1\right)$$\left(\frac{1}{2},1\right)$
Figure S3: The function $f(s)$.
We use the function $f$ to truncate the signed distance function $s(x)$ in (S7) to $f(\frac{s(x)}{d})$, with $d>0$ as in (S10). Notice that $f(\frac{s(x)}{d})=\pm 1$ on $X$, resp. $Y$; hence $f$ only acts like a distance in the white region in Figure S2. For this truncated distance function, we introduce, for all $x\in\Lambda$,
$q_{\mu}(x)=$ $\mathrm{exp}\Big{(}\mu f(\tfrac{s(x)}{d})\Big{)}>0,$ (S12)
$\mu=$ $\frac{da}{2},$ (S13)
where the constant $a$ comes from the short range condition (6). As usual, $q_{\mu}$ is identified with the corresponding multiplication operator.
To lift $q_{\mu}$ to the Fock space $\mathcal{F}(\ell^{2}(\Lambda))$, it is convenient to introduce the following standard notation for the second quantization functor [74].
Definition 4 (Second quantization functor).
Given a one-body operator $A=(A_{xy})_{x,y\in\Lambda}$, we set
$\operatorname{\mathrm{d}\Gamma}(A)=\sum_{x,y\in\Lambda}A_{xy}a_{x}^{\dagger}a_{y},$
and we set $\Gamma(A)=\exp(\operatorname{\mathrm{d}\Gamma}(A))$.
Note the following special case of this definition: If $A_{xy}=q(x)\delta_{xy}$ is a multiplication operator by a function $q(x)$, then
$\operatorname{\mathrm{d}\Gamma}(q)=\sum_{x,y\in\Lambda}q(x)n_{x}.$
The reason for introducing the second quantization functor is that it allows to state various algebraic properties succinctly. Indeed, we will use the following properties which follow directly from the CAR/CCR [74].
Proposition 5 (Properties of $\operatorname{\mathrm{d}\Gamma}$ and $\Gamma$).
  • (i)
    If $A\leq B$, then $\operatorname{\mathrm{d}\Gamma}(A)\leq\operatorname{\mathrm{d}\Gamma}(B)$. In particular, on each $N$-particle sector, we have $\operatorname{\mathrm{d}\Gamma}(A)\leq\|A\|N$.
  • (ii)
    For any function $q:\Lambda\to\mathbb{C}$ and any $x\in\Lambda$, we have the pull-through formulas
    $\Gamma(q)a^{*}_{x}=q(x)a^{*}_{x}\Gamma(q),\quad a_{x}\Gamma(q)=\Gamma(q)\bar{q}(x)a_{x}.$ (S14)
We can now define the central object of the proof.
Definition 6 (Exponential tilting operator).
Set
$T=$ $\Gamma(q_{\mu})=\exp(\mu\operatorname{\mathrm{d}\Gamma}(f(\tfrac{s}{d}))).$ (S15)
Writing this out explicitly,
$T=\exp\left(\mu\sum_{x\in\Lambda}f\left(\tfrac{s(x)}{d}\right)n_{x}\right).$ (S16)
Observe that $T$ is self-adjoint and invertible; see (S12).
We recall the setup of Theorem 1. In particular, we fix the total particle number to be $N$ and we fix two numbers $0\leq\alpha<\beta\leq 1$. The central idea in the exponential tilting method is to simply write
$P_{\bar{N}_{X}\geq\beta}e^{-\mathrm{i}Ht}P_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha}$
$=$ $P_{\bar{N}_{X}\geq\beta}T^{-1}Te^{-\mathrm{i}Ht}T^{-1}TP_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha},$
which leads to the inequality
$\left\lVert P_{\bar{N}_{X}\geq\beta}e^{-\mathrm{i}Ht}P_{\bar{N}_{Y}\geq 1-\alpha}\right\rVert$
$\leq$ $\left\lVert P_{\bar{N}_{X}\geq\beta}T^{-1}\right\rVert\left\lVert Te^{-\mathrm{i}Ht}T^{-1}\right\rVert\left\lVert TP_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha}\right\rVert.$ (S17)
We will now estimate each term of the right-hand-side separately.
Bound on $P_{\bar{N}_{X}\geq\beta}T^{-1}$ and $TP_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha}$
In this section, we prove the following two bounds.
Lemma 7.
On the $N$-particle sector, we have
$\left\lVert P_{\bar{N}_{X}\geq\beta}T^{-1}\right\rVert\leq$ $e^{\mu(1-2\beta)N},$ (S18)
$\left\lVert TP_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha}\right\rVert\leq$ $e^{\mu(2\alpha-1)N}.$ (S19)
Proof.
By Definition (S11), the function $f(\tfrac{s(x)}{d})$ is a regularized version of the map $x\mapsto\frac{2}{d}s(x)$ in the sense that it coincides with $x\mapsto\frac{2}{d}s(x)$ on $W^{0}$ and continuously becomes constant on $X^{\infty}\cup Y^{\infty}=(W^{0})^{\mathrm{c}}$. Explicitly,
$f\left(\frac{s(x)}{d}\right)=\mathbf{1}_{X^{\infty}}(x)-\mathbf{1}_{Y^{\infty}}(x)+\frac{2}{d}s(x)\mathbf{1}_{W^{0}}(x).$
Hence, by the Definition (S15) of $T$,
$T=$ $\mathrm{exp}\Big{(}\mu(N_{X^{\infty}}-N_{Y^{\infty}})+\frac{2\mu}{d}\mathrm{d}\Gamma(s(x)\mathbf{1}_{W^{0}}(x))\Big{)}.$
We now aim to prove the first estimate (S18). Using that $-s(x)\leq\frac{d}{2}$ for $x\in W^{0}$, we find
$-\frac{2\mu}{d}\mathrm{d}\Gamma(s(x)\mathbf{1}_{W^{0}}(x))=-\frac{2\mu}{d}\sum_{x\in W^{0}}s(x)n_{x}\leq\mu N_{W^{0}}.$
As both sides of this operator inequality commute (both operators are diagonal in the occupation basis), this implies
$T^{-1}=$ $\exp\Bigl{(}\mu(N_{Y^{\infty}}-N_{X^{\infty}})-\frac{2\mu}{d}\mathrm{d}\Gamma(s(x)\mathbf{1}_{W^{0}}(x))\Bigr{)}$
$\leq$ $e^{\mu(N_{Y^{\infty}\cup W^{0}}-N_{X^{\infty}})}.$
Recall from Lemma 3 that $X\subset X^{\infty}$ and $Y\subset Y^{\infty}$. Hence, on the subspace $\mathrm{Ran}(P_{\bar{N}_{X}\geq\beta})$, we have $N_{X^{\infty}}\geq N_{X}\geq\beta N$ and $N_{Y^{\infty}\cup W^{0}}\leq N_{X^{\mathrm{c}}}\leq(1-\beta)N$. Combining these estimates, we deduce that
$\left\lVert T^{-1}P_{\bar{N}_{X}\geq\beta}\right\rVert\leq$ $\left\lVert e^{\mu(N_{Y^{\infty}\cup W^{0}}-N_{X^{\infty}})}P_{\bar{N}_{X}\geq\beta}\right\rVert$
$\leq$ $e^{\mu(1-2\beta)N}.$
Since $P_{\bar{N}_{X}\geq\beta}$ and $T^{-1}$ are self-adjoint, we have $\left\lVert P_{\bar{N}_{X}\geq\beta}T^{-1}\right\rVert=\left\lVert T^{-1}P_{\bar{N}_{X}\geq\beta}\right\rVert$ and so (S18) follows.
The second estimate (S19) is proven in the same way, using that $N_{Y^{\infty}}\geq N_{Y}\geq(1-\alpha)N$ and $N_{X^{\infty}\cup W^{0}}\leq N_{Y^{\mathrm{c}}}\leq\alpha N$ on $\mathrm{Ran}(P_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha})$, together with $s(x)\leq\frac{d}{2}$ for $x\in W^{0}$. ∎
Bound on the deformed propagator
To bound (II), it remains to estimate the norm of the deformed propagator $Te^{-\mathrm{i}Ht}T^{-1}$.
Lemma 8.
Suppose that Assumptions (i)(ii) on the Hamiltonian hold. Then, on the $N$-particle sector, we have, for all $t\in\mathbb{R}$,
$\left\lVert Te^{-\mathrm{i}Ht}T^{-1}\right\rVert\leq e^{aNv|t|}.$ (S20)
Proof.
For any bounded operator $A$, we abbreviate $\tilde{A}=TAT^{-1}$. Since $V(\{n_{x}\})$ commutes with $T$, we have $\tilde{H}=\tilde{H}_{0}+V$ and so
$\tilde{U}_{t}=Te^{-\mathrm{i}Ht}T^{-1}=e^{-\mathrm{i}t\tilde{H}}=e^{-\mathrm{i}t(\tilde{H}_{0}+V)}.$
For a bounded operator $A$, we also denote ${\rm Im}A=\frac{A-A^{\dagger}}{2\mathrm{i}}$, which is always self-adjoint. Given any state $\psi$ in the $N$-particle sector, we compute
$\partial_{t}\left\lVert\tilde{U}_{t}\psi\right\rVert^{2}=$ $2\left\langle\tilde{U}_{t}\psi,({\rm Im}\tilde{H}_{0})\tilde{U}_{t}\psi\right\rangle$
$\leq$ $2\sup\mathrm{spec}\,({\rm Im}\tilde{H}_{0})\left\lVert\tilde{U}_{t}\psi\right\rVert^{2}.$ (S21)
Here, for a self-adjoint operator $A$, $\sup\mathrm{spec}(A)$ refers to the supremum over the spectrum of $A$.
Using Gronwall’s lemma and taking the supremum over normalized $N$-particle states $\psi$, it follows that
$\|\tilde{U}_{t}\|\leq\begin{cases}e^{t\sup\mathrm{spec}\,({\rm Im}\tilde{H}_{0})},\quad&t>0,\\ e^{-t\inf\mathrm{spec}\,({\rm Im}\tilde{H}_{0})},\quad&t<0,\end{cases}$
and so
$\|\tilde{U}_{t}\|\leq e^{|t|\|{\rm Im}\tilde{H}_{0}\|}.$ (S22)
It thus remains to bound $\|{\rm Im}\tilde{H}_{0}\|$. We first calculate ${\rm Im}\tilde{H}_{0}$ by using Proposition 5 (ii) with $q=q_{\mu}$ from (S12). This gives
$\tilde{H}_{0}=$ $TH_{0}T^{-1}$
$=$ $\Gamma(q_{\mu})\left(\sum_{x,y\in\Lambda}J_{xy}a_{x}^{*}a_{y}\right)\Gamma(q_{\mu}^{-1})$
$=$ $\sum_{x,y\in\Lambda}J_{xy}q_{\mu}(x)q_{\mu}^{-1}(y)a_{x}^{*}a_{y}$
$=$ $\sum_{x,y\in\Lambda}J_{xy}\exp(\mu(f(\tfrac{s(x)}{d})-f(\tfrac{s(y)}{d})))a_{x}^{*}a_{y}.$
Since $J_{xy}=\bar{J}_{yx}$, we find
${\rm Im}\tilde{H}_{0}=$ $\operatorname{\mathrm{d}\Gamma}(\tilde{J}),$
for $\textnormal{for }\tilde{J}_{xy}=$ $\frac{1}{i}J_{xy}\sinh(\mu(f(\tfrac{s(x)}{d})-f(\tfrac{s(y)}{d}))).$
By Proposition 5 (i), we have
$\|{\rm Im}\tilde{H}_{0}\|=\|\operatorname{\mathrm{d}\Gamma}(\tilde{J})\|\leq N\|\tilde{J}\|$ (S23)
and so it remains to bound the norm of the deformed hopping matrix, $\|\tilde{J}\|$.
To this end, observe that $f(s)$ satisfies $\left\lvert f(s)-f(s^{\prime})\right\rvert\leq\min(2,2\left\lvert s-s^{\prime}\right\rvert)$ for all $s,s^{\prime}\in\mathbb{R}$ (see (S11)). Using this and the fact that $s(x)$ is $1$-Lipschitz, cf. (S8), we have
$\left\lvert f(\tfrac{s(x)}{d})-f(\tfrac{s(y)}{d})\right\rvert\leq\min(2,\tfrac{2}{d}\left\lvert x-y\right\rvert),\quad x,\,y\in\Lambda.$
Recalling that $\mu=\tfrac{da}{2}$, this implies
$\left\lvert\tilde{J}_{xy}\right\rvert\leq\left\lvert J_{xy}\right\rvert\sinh(a\min\{d,|x-y|\}).$ (S24)
By the Schur test for matrix norms and the short-range Assumption (i),
$\|\tilde{J}\|\leq\max_{x\in\Lambda}\sum_{y\in\Lambda}|J_{xy}|\sinh(a|x-y|)\leq av.$
Combining this estimate with (S22) and (S23) proves the lemma. ∎
We now have all the ingredients in place to prove our main result.
Proof of Theorem 1.
Combining (II), (S18), (S19) and (S20) and recalling that $\mu=\frac{a}{2}d_{XY}$, we find that
$\left\lVert P_{\bar{N}_{X}\geq\beta}U_{t}P_{\bar{N}_{Y^{\mathrm{c}}}\leq\alpha}\right\rVert$
$\leq$ $\exp(2\mu(\beta-\alpha)N)\exp(-|t|vaN)$
$=$ $\mathrm{exp}\big{(}-aN\big{[}(\beta-\alpha)d_{XY}-v|t|\big{]}\big{)},$
which proves (8). ∎